Equal masses of water & liquid of density 2 are mixed together, then the mixture has a density of:


a) 2/3   b) 4/3   c) 3/2   d)3

Grade:Upto college level

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago
Dear Mainak

volume of water is =m/1
volume of liquid is = m/2
after mixing   mass of the mixture = (2m)
after mixing volue of mixture =(m+.5m)=1.5m
so density =2m/1.5m
    = 4/3
Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.
We are all IITians and here to help you in your IIT JEE preparation.
All the best.

Regards,
Askiitians Experts
Badiuddin
averonenu
6 years ago
water of density 4 kg/m3 and ice of density 2 kg/m3 are mixed together. if their masses are equal then the density of mixture is?
Arun
6 years ago
Density should be d = 2*d1*d2/(d1 +d2)
 = 2* 4*2/(4+2)
 = 16/6 = 8/3 kg/m³
 
 
Regards
Arun (askIITians forum expert)